OpenStudy (anonymous):

What is the value of the expression 3/3^-2 A. 27 B. 18 C. 1/2 D. 1/3

OpenStudy (anonymous):

@grray7

OpenStudy (michele_laino):

I think is: \[\frac{ 3 }{ \frac{ 1 }{ 9 } }=\frac{ 3 }{ 1 }\times \frac{ 9 }{ 1 }=27\]

OpenStudy (anonymous):

so um what do those brackets and slashes mean? i'm gonna guess 27 do to the equal sign

OpenStudy (michele_laino):

because you have to multiply the first fraction by the inverse of the second fraction, in order to divide two fractions. The fraction you have to divide each other are: 3/1 and 1/9.

OpenStudy (anonymous):

Sorry it's been 10 years since I had to do this, But my nephew thinks I know.

OpenStudy (michele_laino):

So A.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Would you two agree?

OpenStudy (michele_laino):

I think yes!

OpenStudy (anonymous):

I'll take the silence as a yes. He only needs one more mind If I put it here?

OpenStudy (anonymous):

Guess not.

OpenStudy (anonymous):

K thanks anyway

OpenStudy (michele_laino):

@BadMath021402 thanks!